OpenStudy (anonymous):

k is a an integer for which 1/2^1-k, 1/8. What is the value of k? 1/2^1-k must be equal to 1/8 if i read the question right. Missing something simple. Thanx guys.

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 })^{1-k}=\frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

Hi Jonask. Thanx for the help, I appreciate. I think I am giving over the question wrong, Its a GRE question. They give two values 1/2^1-k and then 1/8 in the question. Then they give two possible answers for which they want to know what answer is the biggest, k or -2? Does that explain it better?